Expenditure categories

Million in Birr

1

Personal Consumption Expenditure(C)

4500

2

Gross private domestic investment (I)

850

3

Gov. purchasing of goods and services (G)

1250

4

Export(X)

750

5

Import(M)

1000

6

Net Factor Income received from abroad (NFIRA)

950

7

Net Factor Income paid from abroad (NFIPA)

1050

8

Indirect Business tax (IBT)

450

9

Depreciation(D)

525

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

Based on Above Information calculate

A, Gross Domestic Product

B, Gross National product

C, Net National Income

D, National Income

E, Trade Balance


1
Expert's answer
2022-03-21T12:53:39-0400

a.) Gross Domestic Product (GDP)

"GDP = C + I + G + (X-M)"

"=4500+850 +1250+750-1000"

"=6350"


b.)Gross National Product (GNP)

"GNP=C+I+G+(X-M)+NFIR-NFIPA)"

"=4500+850 +1250+(750-1000)+ (950-1050)"

"=6250"


c.)C) Net National Income (NNP)

"NNP=GND-D"

"NNP=C + I + G + (X-M)-D"

"=4500+850 +1250+750-1000- 525"

"NNP=6350-525"

"= 5,825"


d.)National Income (NI)

"NI=NNP-IBT"

"NNP=6350-525-450"

"=5825-450"

"=53375"


e.) Trade Balance (NX)

"Nx=export -import"

"=750-1000"

"=-250"